OpenStudy (anonymous):

Convert 3pi/5 to degrees

OpenStudy (jdoe0001):

\(\bf degrees = \cfrac{radians \cdot 180}{\pi}\)

OpenStudy (jdoe0001):

\(\bf degrees = \cfrac{radians \cdot 180}{\pi}\implies \cfrac{\frac{3\pi}{5}\cdot 180}{\pi} \)

OpenStudy (jdoe0001):

\(\bf \cfrac{3\pi}{5}\) is a radian unit, thus